# FastRAG: Retrieval Augmented Generation for Semi-structured Data

### v1

@@ III-D3 Text Search (Text)
This strategy employs a slightly different prompting approach. The LLM is given examples of text search features (e.g., from the documentation), typically using regular expressions, and is then instructed to generate a GQL statement that exclusively utilizes text search features to address the input textual query. After the statement is executed the results are interpreted by the LLM.
@@ III-D4 Combined Querying (Combined)
This strategy involves executing both the KG querying and text search prompts in parallel for a given input query. The results from each method are directly returned without LLM interpretation. These results are then provided to the LLM to synthesize the final answer based on the combined context.
@@ III-D5 Hybrid Querying (Hybrid)
This strategy merges the capabilities of both KG querying and text search. The prompt provided to the LLM includes both the KG schema and text search examples, instructing the LLM to generate a GQL statement that can leverage any relevant features from both methods to answer the input query.
IV Evaluation
We designed an experimental setup to evaluate the performance of FastRAG. Our implementation utilized the OpenAI GPT-4o model as the LLM, Neo4j as the graph database to store and query the KG using Cypher GQL, and the Langchain library for LLM prompting.
The evaluation was conducted using two different data sources that represent typical network data: logs and configurations. For the logs, we used a single file containing 2000 lines of OpenStack logs [23], resulting in 1307 chunks of $\sim 1000$ tokens each. For the configurations, we used 13 files of Cisco device configurations, totaling 2100 lines and divided into 19 chunks of $\sim 1000$ tokens each. These datasets were selected to create a manageable use case that allowed for manual verification of the results.
@@ IV-A1 Step1 schema and script learning
When using FastRAG, determining the optimal number of sample chunks for each dataset is crucial, as it directly impacts the quality of schema and script learning. If too few samples are used, the system may overlook important entity types, leading to incomplete parsing. On the other hand, using too many samples can result in overly complex schemas with unnecessary entity types. Additionally, we observed that the optimal sample size for Step 1 does not necessarily translate to Step 2, as the keywords extracted in Step 1 might not fully align with the relevant keywords for each section in Step 2. Consequently, finding the best number of sample chunks for both steps independently is essential.
Samples size is a hyperparameter in our design, requiring experimentation with different sizes by varying $n_{c}$ and $n_{t}$ to identify the sample size that yields the best information extraction. To streamline this process we defined a coverage metric. This metric measures the ratio of lines in the source documents that are included in the input_data of all extracted objects.
To determine the optimal number of chunks for Step 1, we varied the sample size from 1 to M by adjusting $n_{c}$ and $n_{t}$ to achieve the desired number of chunks. We then ran the schema extraction process, generated the corresponding parsing function, and calculated the coverage ratio by executing the function on the source data. Figure 6 reports the results of this procedure. The coverage ratio reaches 1 with just one sample chunk; however, this is because the schema produced is overly simplistic, identifying only one entity type. Therefore, it is also important to consider the completeness and relevance of the extracted schema by examining the number of identified entity types. The goal is to select the most comprehensive schema that also leads to (near) $100\%$ coverage ratio. Another observation is that the LLM tends to add more entity types when more chunks are used, possibly due to the instruction to avoid too many properties in a single entity type. However, we found that overlapping or superfluous entity types are not a significant concern. During script learning, the LLM often fails to find efficient ways to parse these extra types, resulting in empty sections that are ignored in Step 2. This occurs because the LLM prioritizes achieving a 100% coverage ratio, sometimes at the expense of extracting all entity types. To highlight this, Figure 6 differentiates between the number of entity types identified and the number of entity types extracted with the generated parsing function. Based on these observations, we determined the optimal number of sample chunks for each dataset: 4 chunks out of 1307 for the logs dataset and 4 chunks out of 19 for the configuration dataset.
@@
Figure 6: Effect of samples size on extraction (Step 1)
@@
To assess the efficiency of using the LLM in our method, we measured the volume of input and output characters exchanged with the LLM in both schema and script learning prompts, shown in Figure 7. The input character count for both prompts is similar, particularly in the first half of the chart. However, output characters, which weight more in the model’s pricing and are limited to 4000 tokens per output, show different trends. Script learning tends to use fewer output characters, while schema learning’s output increases significantly with larger sample sizes. Nonetheless, with the sample size fixed to $4$ for both datasets, the output character count for schema learning remains within a reasonable range, indicating that the selected sample size is a cost-effective choice.
@@
Figure 7: Effect of samples size on input and output size (Step 1)
We also measured the average latency and total time of the prompts, depicted in Figure 8. These results indicate that the FastRAG processes the entire source data quickly compared to submitting all the data through the LLM. Notably, no prompt call required a retry for any sample size. Schema extraction typically takes longer than code generation, as schemas can be more complex and detailed. However, this process only occurs once at the beginning and remains unchanged unless significant changes, such as new services or applications, are introduced. Similar to character counts, the highest latency and total time values were associated with larger sample sizes, which are beyond the samples size we selected. Only a few requests required retries, mainly in cases where the schema was invalid due to errors like incorrect property types or missing required properties. An unusual spike in the average latency during script learning for the configuration dataset likely resulted from the LLM quickly producing an overly complicated schema that required more time or retries during script learning. This likely led to an imperfect parsing function, as indicated by the coverage ratio of $97\%$.
@@ IV-A2 Step 2 script learning
The process of determining the number of sample chunks for Step 2 script learning is similar to that of Step 1, but it is somewhat simpler. In Step 2, a parsing function is generated for each section, and no schema extraction is involved. As a result, the focus is solely on identifying the number of chunks needed to achieve a coverage ratio close to $100\%$. For simplicity, we fixed the same number of sample chunks for all sections, even though the optimal number might vary between sections. However, the results demonstrate that the number of sample chunks does not need to be precisely accurate; achieving near $100\%$ coverage is possible even with one or two additional chunks beyond what is strictly necessary.
Table I presents the performance metrics and parameters in Step 2 for both datasets. It is worth noting that the prompts generally required more retries than in Step 1. This increase in retries is due to the need to extract detailed properties for each section, which demands a higher level of precision and accuracy from the LLM.
@@ IV-B Question Answering
We evaluate the retrieval effectiveness of FastRAG through question answering. We start by randomly selecting evaluation chunks in Step 1 that were different from those used as sample chunks. We then employed two different LLMs, GPT-4o-mini and Claude-3.5-Sonnet, to generate questions and typical answers based on these evaluation chunks. The questions were submitted to FastRAG and we manually checked the accuracy of the answers. This manual verification is feasible given the relatively small size of the datasets and the manageable number of questions, especially after gaining familiarity with the datasets.
@@
We categorized each answer with one of the following labels: (-) for incorrect answers, (+) for correct answers that do not provide additional information, and (++) for correct answers that include detailed information.
We evaluated FastRAG with the four proposed retrieval methods—Graph, Text, Combined, and Hybrid—on the two studied datasets, logs (Table II) and configurations (Table III).
@@
For the logs dataset, Graph querying retrieval struggled with questions requiring a broader contextual understanding, such as identifying types of logs or understanding the function of nova-compute.log entries. This method often failed to retrieve comprehensive information, especially when abstract or holistic answers were needed. Text retrieval performed better in retrieving specific details directly from the data, such as typical entries in nova-api.log. However, it fell short when exhaustive details were required, as seen in queries about common HTTP methods in the API logs. Combined and Hybrid retrieval outperformed both Graph and Text methods by aggregating information from multiple sources, providing context-rich and detailed answers. Combined retrieval excelled, particularly in identifying the most common HTTP methods and providing a complete explanation of log functions. In certain cases, guiding Text retrieval with query reformulations, such as broadening the search criteria, improved accuracy.
@@
In the configurations dataset, Graph retrieval performed better, particularly in identifying devices and extracting specific configuration details like IP addresses and route-map values. This highlights its effectiveness in relational and accurate data contexts. However, it showed limitations in queries about access lists, possibly due to incomplete data extraction. Text retrieval, meanwhile, frequently produced incomplete results, particularly for more complex queries such as identifying IP addresses or autonomous systems. Once again, Combined retrieval outperformed both individual models, providing more complete answers for queries like listing prefix-lists, thanks to its ability to leverage both text and graph-based methods. Hybrid retrieval, though effective in most cases, occasionally struggled with certain queries, such as access-list matches, pointing to a sensitivity in query formulation.
@@
Overall, combining both graph-based and text-based retrieval provides the most effective solution for question answering tasks in network data, allowing for both precision and completeness. However, while Hybrid retrieval currently underperforms compared to the Combined approach, it shows promising potential for improvement through techniques such as prompt refinement and the use of few-shot examples.
@@
What was the response status for the last GET request to ‘/v2/54fadb41…/servers/detail‘ + + + +
@@
What information can be found in the metadata for the instance logs + - + +
@@
What autonomous systems (AS) are mentioned in the route maps + - + +
What is the most common metric value set in the route maps + + + +
What community value is assigned in the route-map ‘as2_to_as1‘ + - + +
@@ IV-C Cost and Speed Improvement
We conducted an experiment to compare the time and cost of FastRAG with those of the GraphRAG system discussed in Section II, through the processing of logs and configuration data.
To maintain a manageable experiment size while still having enough source data for comparison, we adapted the original datasets as follows: the logs dataset was reduced from 2000 to 500 lines, resulting in 78 chunks (of $\sim 1000$ tokens) for GraphRAG and 319 chunks (of $\sim 1000$ tokens) for FastRAG (Step 1). The configuration dataset was replaced with a larger dataset [24], comprising five configuration files, totaling 4,400 lines, which yielded 67 chunks of $\sim 1000$ tokens for both GraphRAG and FastRAG (Step 1).
To ensure similar conditions for the comparison, we adjusted the parameters of GraphRAG as follows: chunk size was set to 1000 tokens, chunk overlap was set to 50 tokens ($\sim 5$ lines), and the number of concurrent requests was limited to one.
We measured the total time and cost [^2] Cost measured from the API usage provided on the OpenAI dashboard. required to run both RAG systems until the KG creation. For FastRAG, the measured time included the $(n_{c},n_{t})$ hyper-parameters search for the fixed number of sample chunks for each dataset. The procedure of determining the optimal sample size through trial and coverage was not included in the total time. The best sample chunk sizes were fixed as follows. For the configurations dataset: 6 sample chunks in Step 1 (achieving $98.7\%$ coverage) and 4 sample chunks in Step 2 (achieving $94.6\%$ coverage). For the logs dataset: 4 sample chunks in Step 1 (achieving $100\%$ coverage) and 4 sample chunks in Step 2 (achieving $95\%$ coverage).
Table IV presents the results for both evaluated systems. As expected, the difference in cost and time between FastRAG and GraphRAG is significant, even with these relatively small datasets. FastRAG’s improvement is evident, particularly when considering that the generated parsing functions do not need to be regenerated as long as the source data contains the same type of information. These functions can also be updated periodically to refine the information extraction process.
Note that GraphRAG uses the LLM to generate entity descriptions and community reports during the information extraction process, which are essential for the question-answering. Additionally, while we limited GraphRAG to one concurrent LLM request for fairness in this comparison, in practical scenarios, the processing time could be reduced by utilizing parallel requests. This, however, does not reduce the cost.
@@ V Conclusion
FastRAG is a cost-effective and time-efficient method for processing semi-structured network data, demonstrated on network logs and configurations. However, its reliance on sample selection introduces a risk of missing small portions of information ($<5\%$ over the reported results). Additionally, as an LLM-powered system, it introduces variability, requiring continuous verification to ensure accuracy. The current design also lacks entity relationship extraction, which could improve retrieval at a slightly higher cost.
Despite these limitations, FastRAG offers promising performance and resource savings for processing large and frequently changing data, particularly given the rapidly improving LLM capabilities. Interestingly, at the time this study was completed, OpenAI announced a new version of GPT-4o, which reliably adheres to developer-supplied JSON schemas [25], and a new series of models designed to spend more time on reasoning before generating their answers [26].
@@
Any mention of commercial products or reference to commercial organizations is for information only; it does not imply recommendation or endorsement by NIST, nor does it imply that the products mentioned are necessarily the best available for the purpose.

### v2

@@ III-D3 Text Search (Text)
The LLM is given examples of text search features (e.g., from the documentation) and instructed to generate a query that exclusively utilizes text search features to respond to user input. The query is then executed, and the results are interpreted by the LLM.
@@ III-D4 Combined Querying (Combined)
This strategy involves running both the KG query and the text search prompts in parallel for the input of the user. The raw results of each method are then provided to the LLM to synthesize the final answer.
@@ III-D5 Hybrid Querying (Hybrid)
This strategy merges the features of KG querying and text search on the same prompt. The prompt provided to the LLM includes both the KG schema and text search examples and instructs the LLM to generate a query that can use any relevant feature(s) to answer the user input.
IV Implementation and Evaluation
To implement FastRAG, we used the OpenAI GPT-4o model as LLM, Neo4j as the graph database to store and query the KG using Cypher and the Langchain library.
The evaluation used two different data sources that represent typical network data: logs and configurations. For the logs, we used a single file containing 2000 lines of OpenStack logs [21], resulting in 1307 chunks of $\sim 1000$ tokens each. For the configurations, we used 13 Cisco device configuration files, totaling 2100 lines, and divided them into 19 chunks of $\sim 1000$ tokens each.
The sizes were intentionally kept small due to their repetitive syntax and semantics. The heuristic discussed below for determining optimal samples generalizes to larger datasets, making larger size unnecessary for performance evaluation. This approach also allowed for a focused understanding of the data’s content to evaluate question answering.
@@ IV-A1 Step 1 schema and script learning
Determining the optimal number of sample chunks for each dataset is crucial, as it directly affects the quality of schema and script learning. If too few samples are used, the system may overlook important entity types, leading to incomplete parsing. However, too many samples can result in overly complex schemas with unnecessary entity types. Furthermore, we observed that the optimal sample size for Step 1 does not necessarily translate to Step 2, as the keywords extracted in Step 1 might not fully align with the relevant keywords for each section in Step 2. Consequently, it is necessary to find the best number of sample chunks for each step independently.
The size of the samples is a hyperparameter that requires experimentation with different sizes by varying $n_{c}$ and $n_{t}$ to identify the sample size that produces the best extraction of information. To streamline this process, we defined a coverage metric. This metric measures the ratio of lines in the source documents that are included in the input_data of all extracted objects.
To determine the optimal number of chunks for Step 1, we varied the sample size from 1 to M by adjusting $n_{c}$ and $n_{t}$ to achieve the desired number. We then ran the schema extraction process, generated the corresponding parsing function, and calculated the coverage ratio by executing the function on the source data. Figure 4 reports the results of this procedure. The coverage ratio reaches 1 with just one sample chunk; however, the schema produced is overly simplistic, identifying only one entity type. Therefore, it is also important to consider the completeness and relevance of the extracted schema by examining the number of identified entity types. The goal is to select the most comprehensive schema that leads (near) to $100\%$ coverage ratio. Another observation is that the LLM tends to add more entity types when more chunks are used, possibly due to the instruction to avoid too many properties in a single entity type. However, we found that overlapping or superfluous entity types are not a significant concern. During script learning, the LLM often fails to find efficient ways to parse these extra types, resulting in empty sections that are ignored in Step 2. This occurs because the LLM prioritizes achieving a 100% coverage ratio, sometimes at the expense of extracting all entity types. To highlight this, Figure 4 differentiates between the number of entity types identified and the number of entity types extracted with the generated parsing function. Based on these observations, we determined the optimal number of sample chunks for each dataset: 4 chunks out of 1307 for the logs dataset and 4 chunks out of 19 for the configuration dataset.
@@
Figure 4: Effect of sample size on extraction (Step 1)
@@
To assess the efficiency of using the LLM in our method, we measured the volume of input and output characters exchanged with the LLM in both the schema and script learning prompts, shown in Figure 5. The input character count for both prompts is similar, particularly in the first half of the chart. However, output characters, which weigh more in the model’s pricing and are limited to 4000 tokens per output, show a different trend. Script learning tends to use fewer output characters, whereas schema learning’s output increases significantly with larger sample sizes. However, with the sample size fixed to $4$ for both datasets, the output character count for schema learning remains within a reasonable range, indicating that the selected sample size is a cost-effective choice.
@@
Figure 5: Effect of sample size on input and output size (Step 1)
We also measured the prompt’s average latency and total time, depicted in Figure 6. These results indicate that FastRAG processes the source data more quickly than submitting all data through the LLM. In particular, no prompt call required a retry for any sample size. Schema extraction typically takes longer than code generation, as schemas can be more complex and detailed. However, this process only occurs once at the beginning and remains unchanged unless significant changes, such as new services or applications, are introduced in the source data. Similarly to character counts, the highest latency and total time values were associated with larger sample sizes, which are beyond the size of the samples we selected. Only a few requests required retries, mainly in cases where the schema was invalid due to errors like incorrect property types or missing required properties. An unusual spike in the average latency during script learning for the configuration dataset likely resulted from the LLM quickly producing an overly complicated schema that required more time or retries during script learning. This likely led to an imperfect parsing function, as indicated by the coverage ratio of $97\%$.
@@ IV-A2 Step 2 script learning
The process of determining the number of sample chunks for Step 2 script learning is similar to that of Step 1, but somewhat simpler. In Step 2, a parsing function is generated for each section, and no schema extraction is involved. As a result, the focus is solely on identifying the number of chunks needed to achieve a coverage ratio close to $100\%$. For simplicity, we fixed the same number of sample chunks for all sections, although the optimal number might vary between sections. However, the results demonstrate that the number of sample chunks does not need to be precisely accurate; achieving near $100\%$ coverage is possible even with one or two additional chunks beyond what is strictly necessary.
Table I presents the performance metrics and parameters in Step 2 for both datasets. Note that the prompts generally required more retries than in Step 1. This increase in retries is due to the need to extract detailed properties for each section, which demands a higher level of precision and accuracy in the LLM-generated code.
@@ IV-B Question-Answering
We evaluate the retrieval effectiveness of FastRAG through question-answering. To generate relevant questions, we randomly selected evaluation chunks in Step 1 that were different from those used as sample chunks. We then used two different LLMs, namely GPT-4o-mini and Claude-3.5-Sonnet, to generate questions and answers based on these evaluation chunks. The questions were submitted to FastRAG, and we manually checked the accuracy of the answers. This manual verification is feasible given the relatively small size of the datasets and the manageable number of questions.
@@
We submitted the questions to FastRAG with each of the four proposed retrieval methods —Graph, Text, Combined, and Hybrid. We categorized each answer with one of the following labels: (-) for incorrect answers (including hallucinations), (+) for correct answers without any additional information, and (++) for correct answers that include correct additional information.
@@
For the logs dataset (Table II), Graph querying retrieval struggled with questions requiring a broader contextual understanding, such as identifying types of logs or understanding the function of nova-compute.log entries. This method often failed to retrieve comprehensive information, especially when abstract or holistic answers were needed. Text retrieval performed better in retrieving specific details directly from the data, such as typical entries in nova-api.log. However, it fell short when exhaustive information was required, as seen in queries about common HTTP methods in the API logs. Combined and Hybrid retrieval outperformed Graph and Text methods by aggregating information from multiple sources, providing context-rich and detailed answers. Combined retrieval excelled, particularly in identifying the most common HTTP methods and providing a complete explanation of log functions. In some instances, guiding Text retrieval with query reformulations, such as broadening the search criteria, improves accuracy.
@@
In the configurations dataset (Table III), Graph retrieval performed better, particularly in identifying devices and extracting specific configuration details such as IP addresses and route map values. This highlights its effectiveness in relational and accurate data contexts. However, it showed limitations in queries about access lists, possibly due to incomplete data extraction. Text retrieval, meanwhile, frequently produced incomplete results, particularly for queries such as identifying IP addresses or autonomous systems. Once again, Combined retrieval outperformed Graph and Text, providing more complete answers for queries like listing prefix lists, thanks to its ability to leverage both text search and graph-based methods. Hybrid retrieval, though effective in most cases, occasionally struggled with specific queries, such as access-list matches, pointing to a sensitivity in query formulation.
@@
Overall, combining graph-based and text-based retrieval provides the most effective solution for question-answering tasks in network data, allowing for both precision and completeness. However, while Hybrid retrieval currently underperforms compared to the Combined approach, it shows promising potential for improvement through techniques such as prompt refinement and few-shot examples.
@@ IV-C Cost and Speed Improvement
We compared the time and cost of FastRAG with those of GraphRAG discussed in Section II, through the processing of logs and configuration data.
To maintain a manageable experiment size while still having enough source data for comparison, we adapted the original datasets as follows: the logs dataset was reduced from 2000 to 500 lines, resulting in 78 chunks (of $\sim 1000$ tokens) for GraphRAG and 319 chunks (of $\sim 1000$ tokens) for FastRAG (Step 1). The configuration dataset was replaced with a larger dataset [22], comprising five configuration files totaling 4,400 lines, which yielded 67 chunks of $\sim 1000$ tokens for both GraphRAG and FastRAG (Step 1).
To ensure similar conditions for comparison, we adjusted the parameters of GraphRAG as follows: the chunk size was set to 1000 tokens, the chunk overlap was set to 50 tokens ($\sim 5$ lines), and the number of concurrent requests was limited to one  [^2] GraphRAG improves latency by supporting parallel requests to the LLM, an engineering feature that FastRAG could also support but is not implemented in the current prototype.
We measured the total time and cost [^3] Cost measured from the API usage provided on the OpenAI dashboard. required to run both RAG systems until the KG creation. For FastRAG, the measured time included the $(n_{c},n_{t})$ hyperparameter search for the fixed number of sample chunks for each dataset. The procedure of determining the optimal sample size through trial and coverage was not included in the total time. The best sample chunk sizes were fixed as follows. For the configurations dataset: 6 sample chunks in Step 1 (achieving $98.7\%$ coverage) and 4 sample chunks in Step 2 (achieving $94.6\%$ coverage). For the logs dataset: 4 sample chunks in Step 1 (achieving $100\%$ coverage) and 4 sample chunks in Step 2 (achieving $95\%$ coverage).
Table IV presents the results for both evaluated systems. As expected, the difference in cost and time between FastRAG and GraphRAG is significant, even with these relatively small datasets. FastRAG’s improvement is evident, particularly when considering that the generated parsing functions do not need to be regenerated, as long as the source data contains the same type of information. These functions could also be periodically updated to refine the information extraction.
Note that GraphRAG uses LLMs to generate entity descriptions and community reports, which are essential for question-answering. Additionally, while we limited GraphRAG to one concurrent LLM request for fairness, in practical scenarios, the processing time could be reduced by using parallel requests. However, this does not reduce the cost.
@@ V Conclusion
FastRAG is a cost-effective and time-efficient method for processing semi-structured network data, demonstrated on network logs and configurations. Its reliance on sample selection introduces the risk of missing small portions of information ($<5\%$ over the reported results). Furthermore, as an LLM-powered system, it introduces variability, which requires continuous verification to ensure accuracy. The current design also lacks extraction of entity relationships, which could improve information retrieval at a slightly higher cost.
Despite its limitations, FastRAG offers promising performance and resource savings for processing large and frequently changing data, particularly given the rapidly improving LLM capabilities. At the time this study was completed, OpenAI announced a new version of GPT-4o, which reliably adheres to developer-supplied JSON schemas and a new series of models designed to spend more time reasoning before generating their responses.
@@
Any mention of commercial products or references to commercial organizations is for information only; it does not imply NIST recommendation or endorsement, nor does it imply that the products mentioned are necessarily the best available for that purpose.